if pd vs p (where p denotes pressure in atm and d denotes density in g/l) is plotted for He gas (assume ideal) at a particular temprature if (d(pd)/dp)=5,p=8.21 atm .then temprature will be?

Grade:11

1 Answers

Pramod J AskiitiansExpert-IIT-B
13 years ago

Dear student,

PV = mRT/M

=> d = m/V = PM/RT

pd = (P^2)M/RT

d(pd)/dp = 2PM/RT

at P = 8.21 atm d(pd)/dp = 2PM/RT = 5

=> T = 2*8.21*10^5*4/8.314*10^3 *5 = 157.99K

hope this explains.